The following information is extracted from Sheldon Corporation's accounting records at the beginning of 2019.

Accounts Receivable $63000

Allowance for Doubtful Accounts 1400 (credit)

During 2019, sales of credit amounted to $575000 collections on outstanding receivables were $557400 and $2600 of receivables were written off as uncollectible. On December 31, 2019, Sheldon estimates its bad debt to be 4% of the outstanding accounts receivable.

Instructions:

a. Determine the balance in Accounts Receivable at December 31, 2019

b. Prepare the adjusting entry to record Sheldon's bad debt expense for 2019.

c. Prepare the Accounts Receivable section of Shelton's balance sheet at December 31, 2019

d. Compute Shelton's Accounts Receivable turnover and average days collection of receivables
Answer #1

a) Balance in Account receivable = 63000+575000-557400-2600 = 78000

b) Adjusting entry = (78000*4%)-(1400-2600) = 4320

c) Balance sheet

Current assets
Account receivable 78000
Less: Allowance for doubtful accounts -3120 74880

d) Account receivable turnover = 575000/68240 = 8.43

Average days collection = 365/8.43 = 43.30 Days